# On Topology of Three-dimensional Continua

### v1

@@ 1.1 Motivations from solid modeling and CAD
1.2 Motivations from integrating CAD and CAE (Computer Aided Engineering)
@@ 1.4 Contributions of this work
    Based on Definition 4.2 and the answer to (Q-2), can we uniquely represent each 3D continuum by a set of surfaces that are free of overlaps, self-intersections, and proper pairwise intersections?
@@ 4 A surface is a piecewise-smooth orientable compact 2-manifold
Let $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ be two surfaces or surfaces with boundary; they are disjoint if $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. Otherwise a maximal connected component $\gamma$ of $\mathcal{S}_{1}\cap\mathcal{S}_{2}$, which could be an isolated point, a curve, or a surface (with boundary), is called an improper intersection of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ if $(\mathcal{S}_{2}\cap\mathcal{N}(\gamma))\backslash\gamma$ is contained in a single component of $\mathcal{N}(\gamma)\backslash\mathcal{S}_{1}$ in a sufficiently small open neighborhood $\mathcal{N}(\gamma)\subset\mathbb{R}^{3}$; otherwise $\gamma$ is a proper intersection.
@@
For a surface $\cal S$} whose orientation is indicated by a basis $(\mathbf{v}_{1},\mathbf{v}_{2})$ of the tangent space $T_{p}$, the induced normal vector $\mathbf{n}$ at $p\in{\cal S}$ is the vector induced from the right-hand rule to form a basis $(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{n})$ of $\mathbb{R}^{3}$. The surface $\cal S$} is positively oriented if $\mathbf{n}$ always points from its bounded complement to its unbounded complement; otherwise it is negatively oriented.
@@ 5.1 The good neighborhood of a boundary point
By Definition 2.1 and $\cal Y$} being semianalytic, $\partial{\cal Y}\cap\mathcal{N}(p)$ is defined by a finite number of analytic functions $g_{i}:\mathbb{R}^{3}\rightarrow\mathbb{R}$. Since ${\cal Y}\cap\mathcal{N}(p)$ is regular open, $\partial{\cal Y}\cap\mathcal{N}(p)$ contains neither zero-dimensional features such as isolated points nor one-dimensional features such as a curve segment. By the implicit function theorem, the points satisfying $g_{i}(x,y,z)=0$ can be considered as the graph of some height function $\chi=H_{i}(\xi,\eta)$. Hence $\partial{\cal Y}\cap\overline{\mathcal{N}(p)}$ can only contain the homeomorphic images of $\mathbb{D}^{2}$, which, by Definition 5.1, are generalized disks. Therefore (a) holds.
Because each $g_{i}$ is analytic, so is $H_{i}$. Then $H_{i}$ can be replaced by its Taylor series in terms of the local coordinates $\xi$ and $\eta$. For a sufficiently small ${\cal N}(p)$, $H_{i}$ can be approximated arbitrarily well by a bivariate polynomial. In the local neighborhood of $p$, the graphs of two such bivariate polynomials intersect either solely at $p$ or at a curve segment. Therefore (b) holds.
@@ 5.2 The global topology of singular points
    The 0-skeleton of $K_{1}$ is the disjoint union of its 0-cells, written $K^{(0)}:=\sqcup_{i=1}^{k_{0}}p_{i}$.
@@
    Build the 1-skeleton as an attaching space $K^{(1)}:=C^{1}\bigsqcup\nolimits_{f}K^{(0)}$, where the attaching map has the signature $f:\partial C^{1}\to K^{(0)}$ and $C^{1}:=\sqcup_{j=1}^{k_{1}}\gamma_{j}$.
@@
A 1D CW complex is always homeomorphic to a graph $G=(V,E,\phi_{G})$, where the incidence function $\phi_{G}:E\rightarrow V\times V$ can be deduced directly from the attaching map. When the degree of each vertex in $V$ is either 2 or 1, a 1D CW complex is either a 1-manifold or a 1-manifold with boundary; otherwise it is not a 1-manifold. For a Yin set shown in Figure 1, a connected component of its singular points may be one isolated singular point (subplot (a)), a curve segment (subplot (b)), a 1-manifold (subplot (f)), or some 1D subset that is neither a 1-manifold nor a 1-manifold with boundary (subplot (c,d,e)). All these examples are special cases of
@@
    the non-manifold points of $\partial{\cal Y}$ form a compact 1D CW complex, so do the manifold singular points of $\partial{\cal Y}$.
@@
Lemma 5.11 and its proof explain the names “vertex non-manifold points” and “non-vertex non-manifold points” in Definition 3.4.
@@ 5.3 Folded disks
Consider the Yin set in Figure 1(d,e). For each endpoint $q_{e}$ of the horizontal or vertical yellow lines, $\partial{\cal Y}\cap\overline{\mathcal{N}(q_{e})}$ is uniquely represented as a single folded disk centered at $q_{e}$ with $k=2$ and $t=1$; see Figure 4(b). In particular, $\partial{\cal Y}\cap\overline{\mathcal{N}(q_{e})}$ cannot be considered as two disks glued along a single radius at $q_{e}$: although in Figure 4(b) cutting $\cal D$} along $r_{1}$ yields two disks, $p$ would be on the boundary of each disk instead of its interior. In contrast, for the intersection $q_{X}$ of the horizontal and vertical yellow lines, the representation of $\partial{\cal Y}\cap\overline{\mathcal{N}(q_{X})}$ is far from unique: it can be represented as four disks glued along four radii at $q_{X}$, or two folded disks (with $k=2$ and $t=1$) glued along two radii at $q_{X}$, or a folded disk (with $k=4$ and $t=2$) and a disk glued along two radii at $q_{X}$, etc.
@@ 6 Global topology
The unfolding operation continuously deforms each disk ${\cal D}_{i}$ to $\tilde{\cal D}_{i}$. This can be performed one disk at a time so that no new intersections are introduced and can be confined within an arbitrarily small neighborhood of the radius $r$ so that each $\tilde{\cal D}_{i}$ approximates ${\cal D}_{i}$ sufficiently well. Consequently, the unfolding operation eliminates all non-vertex non-manifold points in ${\cal N}(p)$, except those on $\partial{\cal N}(p)$. To understand (c)(i), consider a vertex non-manifold point $p_{v}$ of the Yin set $\cal Y$} in 1(b) and a good neighborhood ${\cal N}(q)$ of a non-vertex non-manifold point $q$ close to $p_{v}$. For ${\cal N}(q)$ satisfying $p_{v}\in\partial{\cal N}(q)$, we cannot separate $p_{v}$ into multiple points because $p_{v}$ is the single center of a folded disk. Indeed, a folded disk in ${\cal N}(q)$ should be transformed to a disk.
Definition 6.2.
@@ 7 Boundary representation
By Lemma 5.11(b), the set $Q$ of all non-manifold points of $\partial{\cal Y}$ form a 1D CW complex homeomorphic to a graph $G=(V,E,\phi_{G})$, where each vertex $v\in V$ is either an isolated non-manifold point or a vertex non-manifold point of $\partial{\cal Y}$ and $E$ contains the curves as the maximal connected subsets of $Q\setminus V$.
@@
A comprehensive example is shown in Figure 6 to illustrate Definition 7.8.
Lemma 7.9.
@@
Theorem 7.11 (Unique B-rep of Yin sets).
For any connected Yin set ${\cal Y}\neq\emptyset,\mathbb{R}^{3}$, its boundary $\partial{\cal Y}$ can be uniquely decomposed into a finite poset $\mathcal{G}_{\partial\mathcal{Y}}=\{\mathcal{S}_{j}\subset\partial{\cal Y}\}$ of pairwise almost disjoint oriented glued surfaces such that
@@
Corollary 7.13.
For any Yin set $\mathcal{Y}\neq\emptyset,\mathbb{R}^{3}$, its boundary $\partial{\cal Y}$ can be uniquely decomposed into a collection ${\cal G}_{\partial{\cal Y}}=\{\{{\cal S}_{i,j}\}:{\cal Y}_{j}\text{ is a connected component of }{\cal Y}\}$ of posets of pairwise almost disjoint oriented glued surfaces such that
(9) $\mathcal{Y}=\bigcup\nolimits^{\perp\perp}_{j}\mathop{\bigcap}\nolimits_{i}\mathrm{int}(\mathcal{S}_{j,i}).$
@@ 8 Conclusion
Acknowledgments. We acknowledge helpful comments from Shaozhen Cao and Chenhao Ye, graduate students at the school of mathematical sciences in Zhejiang University.

### v2

@@ 1.1 Motivations from solid modeling and CAD
1.2 Motivations from integrating CAD and CAE (Computer-Aided Engineering)
@@ 1.4 Contributions of this work
    Based on different types of intersections in Definition 4.2 and the answer to (Q-2), can we uniquely represent each 3D continuum by a set of surfaces that are free of overlaps, self-intersections, and proper pairwise intersections?
@@ 4 A surface is a piecewise-smooth orientable compact 2-manifold
Let $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ be two surfaces or surfaces with boundary; they are disjoint if $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. Otherwise a maximally connected component $\gamma$ of $\mathcal{S}_{1}\cap\mathcal{S}_{2}$ is called an improper intersection of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ if $(\mathcal{S}_{2}\cap\mathcal{N}(\gamma))\backslash\gamma$ is contained in a single component of $\mathcal{N}(\gamma)\backslash\mathcal{S}_{1}$ in a sufficiently small open neighborhood $\mathcal{N}(\gamma)\subset\mathbb{R}^{3}$; otherwise $\gamma$ is a proper intersection.
@@
For a surface $\cal S$} whose orientation is indicated by a basis $(\mathbf{v}_{1},\mathbf{v}_{2})$ of the tangent space $T_{p}{\cal S}$, we regard $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as vectors in $\mathbb{R}^{3}$ by the embedding of $T_{p}{\cal S}$ in $T_{p}\mathbb{R}^{3}$ and define the induced normal vector $\mathbf{n}$ at $p\in{\cal S}$ as the vector induced from the right-hand rule to form a basis $(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{n})$ of $\mathbb{R}^{3}$. The surface $\cal S$} is positively oriented if $\mathbf{n}$ always points from its bounded complement to its unbounded complement; otherwise it is negatively oriented.
@@ 5.1 The good neighborhood of a boundary point
By Definition 2.1 and $\cal Y$} being semianalytic, $\partial{\cal Y}\cap\mathcal{N}(p)$ is defined by a finite number of analytic functions $g_{i}:\mathbb{R}^{3}\rightarrow\mathbb{R}$, each of which can be replaced by its Taylor series expanded at $p$. Thus each $g_{i}$ can be approximated arbitrarily well, within a sufficiently small open neighborhood $\mathcal{N}(p)$, by a linear function of the form $\tilde{g}_{i}(x,y,z)=a_{0}+a_{1}x+a_{2}y+a_{3}z$. Hence the points satisfying $g_{i}(x,y,z)=0$ can be regarded as the graph of some linear height function $\chi=H_{i}(\xi,\eta)$. Therefore, $\partial{\cal Y}\cap\overline{\mathcal{N}(p)}$ can only contain the homeomorphic images of $\mathbb{D}^{2}$, which, by Definition 5.1, are generalized disks. Therefore (a) holds.
The intersection of any two distinct generalized disks in (a) contains no 2D components; otherwise it would contradict Definition 2.2 that any Yin set is regular open. By Definition 5.1, each generalized disk is a surface with boundary and is thus triangularizable. Consequently, ${\cal Y}\cap{\cal N}(p)$ can be approximated arbitrarily well, within a sufficiently small neighborhood ${\cal N}(p)$, by a semianalytic set whose generating functions are multivariate linear polynomials corresponding to triangles in the triangular mesh. The intersection of two such triangular meshes is either $p$ itself or a finite number of curves intersecting only at $p$. Therefore (b) holds.
@@ 5.2 The global topology of singular points
    The 0-skeleton of $K_{1}$ is the disjoint union of its 0-cells, written $K^{(0)}:=\sqcup_{i=1}^{k_{0}}p_{i}$;
@@
    Build the 1-skeleton as an attaching space $K^{(1)}:=C^{1}\bigsqcup\nolimits_{f}K^{(0)}$, where the attaching map has the signature $f:\partial C^{1}\to K^{(0)}$ and $C^{1}:=\sqcup_{j=1}^{k_{1}}\gamma_{j}$;
@@
A 1D CW complex is always homeomorphic to a graph $G=(V,E,\phi_{G})$, where the incidence function $\phi_{G}:E\rightarrow V\times V$ can be deduced directly from the attaching map. When the degree of each vertex in $V$ is either 2 or 1, a 1D CW complex is either a 1-manifold or a 1-manifold with boundary; otherwise it is not a 1-manifold. For a Yin set shown in Figure 1, a connected component of its singular points may be one isolated singular point (subplot (a)), a curve segment (subplot (b)), a 1-manifold (subplot (f)), or some 1D subset that is neither a 1-manifold nor a 1-manifold with boundary (subplots (c,d,e)). All these examples are special cases of
@@
    the non-manifold points of $\partial{\cal Y}$ form a 1D CW complex, so do the manifold singular points of $\partial{\cal Y}$.
@@
The homeomorphism between the 1D CW complex in Lemma 5.11(b) and a graph explains the names “vertex non-manifold points” and “non-vertex non-manifold points” in Definition 3.4.
@@ 5.3 Folded disks
Consider the Yin set in Figure 1(d,e). For each endpoint $q_{e}$ of the horizontal or vertical yellow lines, $\partial{\cal Y}\cap\overline{\mathcal{N}(q_{e})}$ is uniquely represented as a single folded disk centered at $q_{e}$ with $k=2$ and $t=1$; see Figure 4(b). In particular, $\partial{\cal Y}\cap\overline{\mathcal{N}(q_{e})}$ cannot be considered as two disks glued along a single radius at $q_{e}$: although in Figure 4(b) cutting $\cal D$} along $r_{1}$ yields two disks, but according to Definition 5.1, none of them is a disk at $p$. In contrast, for the intersection $q_{X}$ of the horizontal and vertical yellow lines, the representation of $\partial{\cal Y}\cap\overline{\mathcal{N}(q_{X})}$ is far from unique: it can be represented as four disks glued along four radii at $q_{X}$, or two folded disks (with $k=2$ and $t=1$) glued along two radii at $q_{X}$, or a folded disk (with $k=4$ and $t=2$) and a disk glued along two radii at $q_{X}$, etc.
@@ 6 Global topology
The above unfolding operation is well defined since there always exists a function $f$ that continuously deforms each disk ${\cal D}_{i}$ to $\tilde{\cal D}_{i}$ in $\mathbb{R}^{3}$. Furthermore, the unfolding can be performed one disk at a time so that no new intersections are introduced; it can also be confined within an arbitrarily small neighborhood of the radius $r$ so that the max-norm error of approximating each ${\cal D}_{i}$ with $\tilde{\cal D}_{i}$ is no greater than $\epsilon$. Consequently, the unfolding operation eliminates all non-vertex non-manifold points in ${\cal N}(p)$, except those on $\partial{\cal N}(p)$. To understand (c)(i), consider a vertex non-manifold point $p_{v}$ of the Yin set $\cal Y$} in Figure 1(b) and a good neighborhood ${\cal N}(q)$ of a non-vertex non-manifold point $q$ close to $p_{v}$. For ${\cal N}(q)$ satisfying $p_{v}\in\partial{\cal N}(q)$, we cannot separate $p_{v}$ into multiple points because $p_{v}$ is the single center of a folded disk. Indeed, a folded disk in ${\cal N}(q)$ should be transformed to a disk.
Definition 6.2 (Detaching).
@@ 7 Boundary representation
By Lemma 5.11(b), the set $Q$ of all non-manifold points of $\partial{\cal Y}$ form a 1D CW complex homeomorphic to a graph $G=(V,E,\phi_{G})$, where each vertex $v\in V$ is either an isolated non-manifold point or a vertex non-manifold point of $\partial{\cal Y}$ and $E$ contains the curves as the maximally connected subsets of $Q\setminus V$.
@@
We emphasize that the steps (GS.2–4) are necessary because $P_{GS}$ needs not to be a glued surface; see the example shown in Figure 6 to illustrate Definition 7.8.
Lemma 7.9 (Unique glued-surface decomposition of 3D Yin sets).
@@
Theorem 7.11 (Unique B-rep of connected Yin sets).
The boundary $\partial{\cal Y}$ of any connected Yin set ${\cal Y}\neq\emptyset,\mathbb{R}^{3}$ can be uniquely decomposed into a finite poset $\mathcal{G}_{\partial\mathcal{Y}}=\{\mathcal{S}_{j}\subset\partial{\cal Y}\}$ of pairwise almost disjoint oriented glued surfaces such that
@@
Corollary 7.13 (Unique B-rep of Yin sets).
For any Yin set $\mathcal{Y}\neq\emptyset,\mathbb{R}^{3}$, its boundary $\partial{\cal Y}$ can be uniquely decomposed into a collection
(9) ${\cal G}_{\partial{\cal Y}}=\bigl\{{\cal G}_{i}=\{{\cal S}_{i,j}\subset\partial{\cal Y}_{i}\}:{\cal Y}_{i}\text{ is a connected component of }{\cal Y}\bigr\}$
of posets of pairwise almost disjoint oriented glued surfaces such that
(10) $\mathcal{Y}=\bigcup\nolimits^{\perp\perp}_{i}\mathop{\bigcap}\nolimits_{j}\mathrm{int}(\mathcal{S}_{i,j}),$
where each ${\cal G}_{i}$ satisfies Theorem 7.11.
@@ 8 Conclusion
Acknowledgments. We acknowledge helpful comments from Shaozhen Cao, Junxiang Pan, and Chenhao Ye, graduate students at the school of mathematical sciences in Zhejiang University.